OpenStudy (anonymous):

when completing the square r^2-4r=30 the answer would be r=-3.8 and r=7.8 right

OpenStudy (anonymous):

Yes. But I prefer: \[r = 2 \pm \sqrt{34} \] Because inexact answers make me retch.

OpenStudy (anonymous):

No, it's 4. 1/2(b)^2 B is -4, so divide by 2 you get 2, and squared is 4.

OpenStudy (anonymous):

no i mean the final answer

OpenStudy (anonymous):

Oooh sorry didn't see the 30

OpenStudy (anonymous):

so is my answer right

OpenStudy (anonymous):

\[2 + \sqrt{34} \approx 7.8 \] \[2 - \sqrt{34} \approx - 3.8 \]

OpenStudy (anonymous):

OH OK THANK YOU SO IT IS RIGHT

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

THANK YOU